Greenhab Report – December 23rd

Greenhab Report
David Murray
23 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with door open from 9:56 to 15:37
Working Hour: 14:54
Inside temp at working hour: 19 C
Outside temp during working hours: 4.5 C
Inside temperature high: 25 C
Inside temperature low: 15 C
Inside humidity: 16%
Heating
Functioning nominally,
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
Kale and cabbage are sprouting
Morning Research Observations
N/A

Daily Water Usage for Crops
2 gallons
Time(s) of Watering Crops
15:37
Aquaponics
N/A
Narrative
Today I marked the controls and applied the Moringa oleifera fertilizer to 13 tomato plants. The kale and cabbage are starting to sprout but no sign of the spinach. Space in the greenhab is limited so I have to remove some of the tomatoes.
Support/Supplies Needed
N/A

Operations Report – December 23rd

Crew 185

Operations Report Saturday 12/23/2017

SOL: 5

Name of person filing report: A.Passeron

Non-nominal systems:

– Generator :
The generator had a failure during the night, Fault message reading "Low battery". Jump-start from Shannon’s Shuttle battery at did not solve the issue. Generator had a second failure at rougthly 12:00pm. Second jump-start done and generator turned on at 5:30pm. The generator is currently running.
At 6:20pm, settings are :
+ Oil pressure : 45 psi
+ Speed : 1815 rpm
+ Battery charge : Fluctuating between 8.7V and 9.0V
At 6:40pm, settings are :
+ Oil pressure : 45 psi
+ Speed : 1815 rpm
+ Battery charge : Fluctuating between 7.8V and 8.0V
The charge of the battery is thus plummeting. The battery of the generator is not recharged while generator is on.

– Water heater tests :
I went to the shower and did the following tests.

@8:42am first trial with nominal settings (right dial on the second top black bar) :
hot water turned on all the way for 2 minutes, initial temperature 51F, final temperature 64F.
@9:33am second trial with max temperature (right dial turned to the top black bar, but I can go further) :
hot water turned on all the way for 4 minutes, initial temperature 55F, final temperature 64F.

– Water shortage :
Crew aware of concern about water shortage. We are now using jerrycans for drinking water. We tried to pump as much water as possible from the trailer tanks into the static tank. We use the shallow water of trailer tanks to water the plants of the GreenLab.

Generator (hours run):
Generator turned off at unknown time (failure during the night)
Jump-start done at 8:10am
Generator turned on at 8:10am
Generator turned off at roughtly 12:00pm (second failure during the EVA)
Jump-start done at 5:30pm
Generator turned on at 5:30pm

Solar— SOC
@ 8:00am : 57%
@ 9:20am : 60%
@ 10:00am : 64%
@ 11:00am : In EVA
@ 12:00pm : In EVA
@ 1:00pm : In EVA
@ 2:00pm : 100%
@ 3:00pm : 98%
@ 4:50pm : 90%
@ 5:00pm : 88%
@ 5:30pm : 84%
@ 6:40pm : 82%

Diesel – 78%

Propane – 52%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2.8 Gallons

Water (trailer) – 20 gallons

Water (static) – 280 gallons

Trailer to Static Pump used – Yes

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: Yes

ATVs Used: Yamaha blue Grizzly 350 #3
Oil Added? Yes
ATV Fuel Used:
# Hours the ATVs were used today: 02:00 hours
Notes on ATVs: ATVs were nominal.

Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Spirit rover used: Yes
Hours: 02:00 Hours
Beginning charge: 100%
Ending charge: 60%
Currently charging: Yes

Opportunity rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: Water shortage prevention. Water heater doesn’t seem to work properly.

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

EVA Report – December 23rd

Report: Crew 185 EVA #4 for SOL 5.
Propose of EVA: Collecting geological samples for mobility testing. Then, walking with the EVA suits and discover the terrain around our position.

EVA Report:
(i) The first 40 min were dedicated to the engineering check during which we have filled the static tank with the left unfrozen water of the tanks on the trailer. Then, using a pan, we collected other (frozen) water that can be used for the Green Hab.
(ii) Arrived at the White Moon, we did a short walk around. Because strangers were near the vehicles, we went back, and we collected a few geological samples.
(iii) Then, we stopped about 1 km from the Hab to locate recognition signs to be used for the following two EVAs.
(iv) Back at the Hab, we refilled the ATV 350 and plugged the rover.

EVA #5:
Start time: 1030am
End time: 1230pm -> finished at 2pm

Ilaria Cinelli: Leader
John : Astro 1
Arno: Astro 2
Destination: White Moon (110.8W,38.44N).
Vehicles: 1 ATV (Ilaria), 1 Rover (John and Arno)
CapCom(s): David and Thibo

Written by Commander Ilaria Cinelli

Sol Summary – December 23rd

Crew 185 Sol Summary Report 23 Dec 2017

Sol 5

Beginning of a Navigation Exercise

David Murray

Mission Status: We had slight generator problems but everything is under control and we still have charge.

Sol Activity Summary: The crew ate breakfast and disbanded for individual tasks before meeting back up to do the EVA prep. Ilaria, John, and Arno made up the EVA crew today. They placed objects that the next EVA crew will have to locate in the field based on instructions from CapComm. After the EVA Arno and I participated in John’s medical training.

Look Ahead Plan: Tomorrow we’re planning to do a split EVA to test the X-1 and to locate the items that today’s EVA crew put in the field.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Healthy

EVA: 3 person EVA to use ATV’s and hide objects for a navigation study.

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, EVA Report, Sol Summary

Support Requested: None